I can't figure out how to set this problem up. Your help would be greatly appreciated.
An airplane travels 400 miles against the wind in 4 hours, and makes the return trip with the same wind in 2 hours. Find the speed of the wind.
Thank you.

Answer by Paul(988)

Let the speed of the plane be x
Let the speed of the wind be y
2x+2y=400
y=200-x (subsitution)
4x-4y=400
4x-4(200-x)=400
8x=1200
x=150
y=200-150
y=50
Hence, the speed of the plane is 150mph and the speed of the wind is 50mph.
PAul.